OpenStudy (anonymous):

Am I doing this problem right? 6^((log_36)^4)=16

OpenStudy (anonymous):

6^((log_36)^4)=6 ^ ( log

OpenStudy (anonymous):

what base log is that?

OpenStudy (anonymous):

6^((log_36)^4)=6 ^ ( (log 6^2 )^4 = 2^4 = 16

OpenStudy (anonymous):

Thats what I did, but when I put it into my hmwk (online) then it was wrong...?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

6^((log_36)^4) = 6 ^ [ (log_6 6^2) ^4 ] = 6 ^ [ 2^4] = 6^16

OpenStudy (anonymous):

Oh, so I cancelled too many 6's?

OpenStudy (anonymous):

dunno, did my solution make sense?

OpenStudy (anonymous):

2 821 109 907 456

OpenStudy (anonymous):

um, it did but I'm confused still. It doesn't match very well with what I'm getting, well it is but it's not pretty.

OpenStudy (anonymous):

but thats the right answer, correct ?

OpenStudy (anonymous):

6^16=2.821109907e^12 , not sure gonna try

OpenStudy (anonymous):

not right

OpenStudy (anonymous):

no dont approximate

OpenStudy (anonymous):

2 821 109 907 456

OpenStudy (anonymous):

why no decimal

OpenStudy (anonymous):

theres no need for it

OpenStudy (dumbcow):

what base is the log?

OpenStudy (anonymous):

went to the end, and it showed me the answer is 2

OpenStudy (anonymous):

yeah, you probably should give us the whole question

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

it says to simplify that equation i gave

OpenStudy (dumbcow):

what base is the log?

OpenStudy (anonymous):

I'm confused about that to, i think its 6

OpenStudy (anonymous):

\[6^{\log _{36^{4}}} = ..... .\log _{6^{x}}=\log6^{2\times4}......x=8..........6^{8}\] ?

OpenStudy (anonymous):

no, I dont think the 4 is a power of 36

OpenStudy (anonymous):

at the bottom eqution can u try rewrite

OpenStudy (anonymous):

It's ok, I moved on from that question so you dont need to waste your brain power on this question. Thanks for trying :)

OpenStudy (anonymous):

\[6^{\log _{36}4}\]

OpenStudy (dumbcow):

ahh base is 36 and 4 is not an exponent x = log base 36 (4) => 36^x = 4 => 6^2x = 4 thus x = log base 6 (4) /2 6^x = sqrt(4) = 2

OpenStudy (anonymous):

I didn't understand that other answer, but it went away before i could finish looking. I understand yours a little better.

OpenStudy (anonymous):

\[36=6^{2}\]\[36^{x}=4\]\[6^{2(x)}=4\]\[2x=4\]\[x=\frac{4}{2}\]\[x=2\]